# Quantum-inspired Techniques in Tensor Networks for Industrial Contexts

### v1

@@ 1 Introduction
On the other hand, one of the most promising technologies, although more unknown within the industry, is tensor networks (TN) [14]. Tensor networks are a class of quantum-inspired algorithms and techniques based on mimicking the tensor operations performed by a quantum computer, but executing them on classical computers. By using tensor properties, the execution of such operations can be optimized, especially in cases where the entire quantum state vector is not required, but only properties of it.
Another relevant feature of tensor networks is that they allow the efficient representation of certain families of quantum states, by means of representations such as the matrix product state (MPS)[15, 16], also called tensor train (TT), or projected entangled pair states (PEPS) [16]. In this way, with a reduced amount of memory we can perform quantum calculations and obtain properties of complex systems. These same representations have been relevant to the world of machine learning, since they are capable of compressing models by reducing the memory required without losing a notable amount of precision [17, 18, 19].
Given all these capabilities, tensor networks are highly susceptible to be used in industrial contexts, being able to address highly complex and large problems efficiently. Despite being a relatively mature field at the academic level (initiated in 1971 by Penrose [20]), its study applied to industry is recent and always associated with quantum computing. However, given the great results achieved in recent years by tensor networks, outperforming quantum algorithms executable on current hardware [21, 22], many companies are beginning to take an interest in the study and implementation of this technology.
@@ 2 Introduction to tensor networks
A highly interesting and useful type of tensor network is the already mentioned MPS [15], since it consists of an one-dimensional and decomposed representation of a tensor by means of the product of matrices. We can see an example in Fig. 2 a), where we see the MPS representation of a 5-tensor. This MPS representation is composed of 2 types of indexes: the bond indexes, which will have dimension $b$, and the physical indexes, of dimension $d$, corresponding to the dimensions of the tensor to be represented. The physical indexes are the free indexes, which coincide with the indexes of the tensor to be represented, while the bond indexes are those that connect the tensors of the network to each other and are responsible for accounting for the correlations between the physical indexes. It should be noted that there are other versions of the MPS representation with different boundary conditions (the way to link the end nodes), but the simplest is this one. There is also the possibility that the dimensions of the bond indexes are different, as are those of the physical indexes, for reasons we will mention at the end of this section.
@@
Any tensor can be represented in MPS representation exactly, although not always efficiently, since $b$ can be of the order of $d^{N}$ in certain complex cases and where the correlation is very strong. Therefore, it will be important to consider how $b$ increases with the size of the tensor to be represented. Still, many cases of interest can be represented in MPS representation efficiently. For example, in academic contexts its most widespread use is that of representing quantum states of $N$ particles, that is, state vectors of $d^{N}$ components.
@@
Such forms can be obtained mainly in three ways. The first is by logical construction of the representation, so that the interactions between the different nodes of the tensor network generate the tensor we want. The second will be the variational construction, consisting of starting with a test set of tensors and using techniques such as gradient descent to get it to become the one we want to represent. The third is by singular value decomposition (SVD) [25], iterative, where we perform an SVD for each pair of physical indexes. There is also an optimized method for sparse tensors [26].
@@
Consider an example of obtaining the MPO representation of $n$ nodes of a $N\times M$ matrix. To do this we will have to start by transforming the matrix into a $2n$-tensor, by means of a splitting. The resulting $A^{\prime}$ tensor will have dimensions $\vec{D}$ such that
$N=\prod_{i=0}^{n-1}D_{i},\qquad M=\prod_{i=n}^{2n-1}D_{i}.$ (2)
Now we join the pairs of input and output indexes that we want to be part of the same node of the MPO form and we group them to obtain a tensor $A^{\prime\prime}$ whose new dimensions $d_{i}=D_{2i}D_{2i+1}$ are $d_{i}=D_{2i}D_{2i+1}$. Now we perform the iterative process of searching for the MPS representation of that tensor, to finally split the physical indexes of each node to obtain the original input and output indexes. The whole process can be seen in Fig. 3.
@@
If instead of wanting an exact representation of a tensor, we would like an approximate representation of it in order to have a smaller $b$ and thus save memory, we can take advantage of these two representations. In the variational construction we will simply operate with smaller $b$, while in the iterative SVD we will perform truncated SVD (TSVD) [27], keeping only the $b$ largest singular values, which represent the most significant correlations. We can also approximate a previously obtained MPS representation [28]. The accuracy of this approximation will depend both on the tensor itself that we want to represent and on the mapping we use, since if we put together in the same nodes the most correlated ‘columns’, all their correlations will be within the node itself. Moreover, if we have the most correlated indexes adjacent to each other, we will need a smaller bond dimension, since the intensity of the correlations between two nodes of an MPS representation decays exponentially with the geodesic distance between them, the minimum distance that separates them.
@@ 3.2 Machine learning model compression
where $A$ is a matrix of dimension $N\times M$, $\vec{v}$ and $\vec{c}$ are vectors of dimension $N$ and $\vec{x}$ is a vector of dimension $M$. The input to the layer is the vector $\vec{x}$, while the output will be the vector $\vec{v}$. The layer will be characterized by the $M$ components of the matrix $A$ and the $N$ components of the vector $\vec{c}$, which will be the ones to be trained. Therefore, the layer will have a number of $N(M+1)$ parameters. To reduce the number of parameters we can represent the matrix $A$ as an approximate MPO form with a certain bond dimension $b$ so that we have a smaller number of parameters than in the original dense version. We can do the same with the vector $\vec{c}$ obtaining its approximate MPS form. In this way, we have compressed the model layer. This is favorable both to save memory space and to train the model faster and avoid problems such as overfitting, depending on the representation used and the original model.
@@
There are many compression schemes in addition to the MPS and MPO, such as the PEPS [34, 35] or the MERA [36]. These allow to efficiently perform certain operations on their compressed representations and to apply various optimization methods, such as the Density Matrix Renormalization Group (DMRG) [36, 37].
@@ 3.3 Application of high dimensionality operations
where each kernel $\phi_{j}(\cdot)$ affects the $j$-th component of the vector $\vec{x}$. In this case, we can consider a kernel $j(\cdot)$.
$\phi_{j}(x_{j})=(x_{j},1),$ (5)
so that the global kernel $\phi(\cdot)$ will give us all possible products between the components of the input vector, as a $2^{N}$ component tensor to express correlations in the vector. As we can see, in the standard way this would require an exponential amount of memory and time in the size of the input vector. Furthermore, if we want to apply a matrix to such a kernel we would need a matrix $A$ of dimension $M\times 2^{N}$, which scales prohibitively.
However, by taking advantage of the properties of tensor networks we can use an MPO representation for such a matrix, of the form depicted in Fig. 4, where only $2N$ smaller tensors need to be contracted.
@@
With this methodology we can apply an operation in which we weight elements of type $x_{1}x_{2}x_{3}x_{4}x_{5}$, $x_{1}x_{2}x_{3}x_{4}$, $x_{1}x_{2}x_{3}$, $x_{1}x_{2}x_{2}$ and $x_{1}$, efficiently and with a reduced number of elements.
@@ 3.4 Data compression and denoising
In the field of big data there is a need to work with large volumes of data, which can be very expensive or unfeasible for certain systems. The tensor compression property that we have seen in representations such as MPS allow us to compress [40] and handle [41] large amounts of data. To do this we just need to convert the data into a tensor and compress it. In addition, if we allow ourselves to reduce the precision of its elements or are only interested in its global characteristics, we can use approximate representations of the tensor to save memory space.
@@ 3.5 Combinatorial optimization
Solving most combinatorial optimization problems requires runtimes exponentially increasing with problem size. Given the properties of quantum computation, many can be tackled through it efficiently. A similar approach to quantum lies in the use of tensor networks for combinatorial optimization. There are different types of combinatorial optimization solvers.
The first one is the one developed in [42], running in a black box. This algorithm starts from a cost tensor, $T$, whose subscripts are the components of the solution vector. This tensor will have the cost of all possible states, although it is not necessary to know the tensor explicitly: it is enough to have a function that gives us the cost as a function of its indexes. With this tensor, we will make an approximation of its max volume submatrix, the submatrix containing its largest singular values in product, by means of an MPS representation. From this representation we will obtain, with high probability, the lowest cost element.
@@
The third possibility is the use of imaginary time evolution to a quantum state in uniform superposition of all possible combinations, equivalent to the application of a Hadamard gate to each qubit of a quantum circuit. If from this uniform superposition we damp each combination exponentially to its associated cost, the optimal combination will be the one with the largest amplitude, resulting in a higher measurement probability.
@@
Due to its non-unitareity, at the quantum level this type of operation cannot be performed, but at the classical level it is possible to perform [44]. If we needed to take into account all possible combinations, we would need an unbearable amount of memory, resulting in a brute-force search. However, by combining the ability to compress information and simulate quantum states with the ability to apply non-unitary tensor network operations, we can create such a state for various problems [45, 46, 47]. This can be done in an exact or approximate way, by recompressions, depending on how much memory we want to devote to it.
@@
The main limitation is that for NP-Hard problems, the most complex ones, we will need an exponential amount of memory and time to solve them exactly. By making intelligent adjustments in the encoding and expression of the problem, the creation and contraction of the tensor network and applying various simplifications in the constraints, we can create efficient approximate algorithms from the exact ones, avoiding exponential scaling.
@@
The fourth is the simulation of a physical system represented by a variational tensor network, with a cost function given as a Hamiltonian that we can write in MPS or MPO representation. In this case we will perform variational updates of the state tensor network, using techniques such as gradient descent or DMRG to minimize the energy of the system to find the optimal combination. The effectiveness of these types of techniques depends very much on the problem to which they are applied and they can get stuck in local minima, especially for problems with constraints. In addition, there is the limitation that we must be able to express the cost function in a tensor network representation.

### v2

@@ 1 Introduction
On the other hand, one of the most promising technologies, although less known within industry, is tensor networks (TN) [7]. Tensor networks are a class of quantum-inspired algorithms and techniques based on mimicking the tensor operations performed by a quantum computer, but executing them on classical computers. By using tensor properties, the execution of such operations can be optimized, especially in cases where the entire quantum state vector is not required, but only properties of it.
Another relevant feature of tensor networks is that they allow the efficient representation of certain families of quantum states, by means of representations such as the matrix product state (MPS)[54, 18], also called tensor train (TT), or projected entangled pair states (PEPS) [18]. In this way, with a reduced amount of memory we can perform quantum calculations and obtain properties of complex systems. These same representations have also been relevant to machine learning, since they can compress models by reducing the memory required without necessarily losing a notable amount of precision [57, 70, 78].
Given all these capabilities, tensor networks are relevant to industrial contexts when the chosen tensorization exposes low-rank, local, hierarchical, sparse, or topology-aligned structure. When the required ranks or contraction widths grow rapidly, the same representations may become impractical. Despite being a relatively mature field at the academic level (initiated in 1971 by Penrose [53]), its study applied to industry is more recent and often associated with quantum computing. At the same time, recent tensor-network simulations have also become part of the classical benchmark used to assess quantum-hardware claims [69, 52, 36], which has increased interest in their practical deployment.
@@
1.1 Tensor networks in 2026: scope and current role
As of 2026, tensor networks occupy a broader role than the classical simulation of isolated quantum systems. They are used as compressed representations for quantum many-body states and quantum circuit simulation, as strong classical baselines for quantum-advantage claims, and as quantum-inspired numerical tools in machine learning, optimization, finance, scientific computing, and data compression [36, 71]. In this sense, tensor networks are not only quantum-inspired algorithms but also part of the classical benchmark that any claim of practical quantum advantage must confront.
A further change is the maturation of the surrounding software stack. Contraction planning, GPU execution, and tensor-operation interfaces have become more standardized and more relevant for deployment, which makes implementation details increasingly important in industrial assessments [8]. Across all these roles, the central limitation remains structural: tensor-network methods are effective when the data, model, operator, or feasible set admit controlled ranks and manageable contraction widths. They are therefore not generic black-box cures for high dimensionality or NP-hardness.
@@ 2 Introduction to tensor networks
A highly interesting and useful type of tensor network is the already mentioned MPS [54], since it consists of a one-dimensional decomposed representation of a tensor by means of the product of matrices. We can see an example in Fig. 2 a), where we see the MPS representation of a 5-tensor. This MPS representation is composed of 2 types of indexes: the bond indexes, which will have dimension $b$, and the physical indexes, of dimension $d$, corresponding to the dimensions of the tensor to be represented. The physical indexes are the free indexes, which coincide with the indexes of the tensor to be represented, while the bond indexes are those that connect the tensors of the network to each other and are responsible for accounting for the correlations between the physical indexes. It should be noted that there are other versions of the MPS representation with different boundary conditions (the way to link the end nodes), but the simplest is this one. There is also the possibility that the dimensions of the bond indexes are different, as are those of the physical indexes, for reasons we will mention at the end of this section.
@@
Any tensor admits an exact open-boundary MPS/TT representation, although not always efficiently [54, 51]. The relevant question is how the bond dimensions scale with the tensor size and with the ordering of the indexes.
Proposition 2.1.
Let $X\in\mathbb{F}^{d_{1}\times\cdots\times d_{N}}$, and let $X^{\langle k\rangle}\in\mathbb{F}^{(d_{1}\cdots d_{k})\times(d_{k+1}\cdots d_{N})}$ be the unfolding across the bipartition $(1,\dots,k)\mid(k+1,\dots,N)$. Then there exists an exact open-boundary TT/MPS representation whose minimal bond dimensions satisfy
$r_{k}=\operatorname{rank}\left(X^{\langle k\rangle}\right),\qquad k=1,\dots,N-1.$ (2)
Consequently,
$r_{k}\leq\min\left(d_{1}\cdots d_{k},\ d_{k+1}\cdots d_{N}\right),$ (3)
and if $d_{j}=d$ for all $j$, then
$\max_{k}r_{k}\leq d^{\lfloor N/2\rfloor}.$ (4)
Generic tensors attain these maximal unfolding ranks.
Proof.
Successive SVDs across the cuts $(1,\dots,k)\mid(k+1,\dots,N)$ produce an exact TT/MPS with bond dimensions $\operatorname{rank}(X^{\langle k\rangle})$. Conversely, cutting any exact TT/MPS between nodes $k$ and $k+1$ factorizes $X^{\langle k\rangle}$ through the bond space of dimension $r_{k}$, so $\operatorname{rank}(X^{\langle k\rangle})\leq r_{k}$. Hence the minimal bond dimension on that cut is exactly $\operatorname{rank}(X^{\langle k\rangle})$. ∎
Therefore, exact MPS representations can still be exponentially large, since for uniform physical dimension $d$ one may need $b\geq d^{\lfloor N/2\rfloor}$. Efficient tensor-network methods rely instead on low effective ranks or on controlled truncation. For example, in academic contexts a widespread use is to represent quantum states of $N$ particles, that is, state vectors of $d^{N}$ components.
@@
Such forms can be obtained mainly in three ways. The first is by logical construction of the representation, so that the interactions between the different nodes of the tensor network generate the tensor we want. The second will be the variational construction, consisting of starting with a test set of tensors and using techniques such as gradient descent to get it to become the one we want to represent. The third is by iterative singular value decomposition (SVD) [81, 51], where we reshape the tensor across successive bipartitions and apply SVDs recursively. There is also an optimized method for sparse tensors [38].
@@
Consider a matrix $A\in\mathbb{F}^{P\times Q}$ that we want to represent as an MPO with $n$ nodes. To do this we first split the row and column indexes according to factorizations
$P=\prod_{k=1}^{n}p_{k},\qquad Q=\prod_{k=1}^{n}q_{k}.$ (5)
This identifies the global row and column indexes with tuples
$a\leftrightarrow(a_{1},\dots,a_{n}),\qquad b\leftrightarrow(b_{1},\dots,b_{n}),$ (6)
where $a_{k}\in\{1,\dots,p_{k}\}$ and $b_{k}\in\{1,\dots,q_{k}\}$, and therefore
$A_{a,b}=\mathcal{A}_{a_{1},\dots,a_{n},b_{1},\dots,b_{n}}.$ (7)
To construct the MPO, we permute the indexes to $\widetilde{\mathcal{A}}_{a_{1},b_{1},a_{2},b_{2},\dots,a_{n},b_{n}}$ and group each pair $(a_{k},b_{k})$ into a local index of dimension $d_{k}=p_{k}q_{k}$. We then apply TT-SVD to this reordered tensor and finally split each local index again to recover the two physical indexes per MPO node. In the notation of a $2n$-tensor with dimensions $(D_{0},\dots,D_{2n-1})$, if the first $n$ dimensions correspond to the rows and the last $n$ to the columns, then the correct local dimensions are
$d_{i}=D_{i}D_{n+i},\qquad i=0,\dots,n-1.$ (8)
The formula $d_{i}=D_{2i}D_{2i+1}$ is valid only after explicitly interleaving the dimensions as $(D_{0},D_{n},D_{1},D_{n+1},\dots,D_{n-1},D_{2n-1})$. The whole process can be seen in Fig. 3.
@@
If instead of wanting an exact representation of a tensor, we would like an approximate representation of it in order to have a smaller $b$ and thus save memory, we can take advantage of these two representations. In the variational construction we will simply operate with smaller $b$, while in the iterative SVD we will perform truncated SVD (TSVD) [24], keeping only the largest singular values on each unfolding. We can also approximate a previously obtained MPS representation [51]. In TT-SVD, if the discarded squared singular-value mass at step $k$ is $\epsilon_{k}^{2}$, then the resulting approximation $\widetilde{X}$ satisfies
$\|X-\widetilde{X}\|_{F}^{2}\leq\sum_{k=1}^{N-1}\epsilon_{k}^{2}.$ (9)
Thus, choosing $\epsilon_{k}=\epsilon/\sqrt{N-1}$ yields $\|X-\widetilde{X}\|_{F}\leq\epsilon$, and for fixed TT ranks the construction is quasi-optimal up to a factor $\sqrt{N-1}$ in Frobenius norm. The accuracy of this approximation will depend both on the tensor itself that we want to represent and on the mapping we use, since if we put together in the same nodes the most correlated ‘columns’, all their correlations will be within the node itself. Moreover, bringing strongly correlated indexes close to each other can reduce the bond dimensions needed for a good approximation. For injective homogeneous MPS, correlations of local observables decay exponentially with distance when the transfer matrix has a nonzero spectral gap. More generally, the MPS topology constrains entanglement across cuts through the bond dimensions, which is why the ordering of variables affects the ranks needed for an accurate representation.
@@ 3.1 Simulation of quantum systems
Beyond many-body simulation, tensor networks have become a central tool for assessing claims of quantum advantage. Several recent quantum-hardware experiments have been analyzed by comparing them with optimized tensor-network simulations [69, 52, 36]. This does not remove the relevance of quantum hardware; rather, it raises the standard for advantage claims and clarifies the regimes in which classical contraction methods remain competitive.
@@ 3.2 Machine learning model compression
where $A$ is a matrix of dimension $N\times M$, $\vec{v}$ and $\vec{c}$ are vectors of dimension $N$ and $\vec{x}$ is a vector of dimension $M$. The input to the layer is the vector $\vec{x}$, while the output will be the vector $\vec{v}$. The layer has $NM$ trainable parameters in the matrix $A$ and $N$ trainable parameters in the bias vector $\vec{c}$, for a total of $NM+N=N(M+1)$ parameters before compression. To reduce the number of parameters we can represent the matrix $A$ as an approximate MPO form with a certain bond dimension $b$ so that we have a smaller number of parameters than in the original dense version. We can do the same with the vector $\vec{c}$ obtaining its approximate MPS form. In this way, we have compressed the model layer. This is favorable both to save memory space and to train the model faster and avoid problems such as overfitting, depending on the representation used and the original model.
@@
There are many compression schemes in addition to the MPS and MPO, such as the PEPS [11, 74] or the MERA [58]. These allow certain operations to be performed efficiently on their compressed representations and can be optimized with local variational methods, including DMRG in the MPS/one-dimensional setting and related sweep or variational algorithms for PEPS, MERA and other tensor-network geometries [58, 30].
From 2024 onward, much of the applied focus has shifted from compressing relatively small neural networks to compressing transformer and large language model components. MPO/TT decompositions can be applied to dense or linear layers, attention projections, and feed-forward blocks, while tensorized adapters provide parameter-efficient fine-tuning mechanisms [70, 71, 41, 31].
However, parameter compression alone does not guarantee lower wall-clock latency. Practical acceleration requires training and inference to operate directly on the compressed cores, together with efficient contraction kernels; reconstructing dense matrices removes much of the benefit. For industrial AI, the relevant question is therefore not only the compression ratio, but the full deployment path: whether the compressed cores are used natively during training and inference, whether hardware kernels support the resulting contractions, and whether the accuracy–latency–memory trade-off improves over quantization, pruning, distillation, LoRA-type adapters, and conventional low-rank decompositions [70, 41, 31, 43].
@@ 3.3 Application of high dimensionality operations
where each local feature map $\phi_{j}(\cdot)$ acts on the $j$-th component of the vector $\vec{x}$. In this case, we can consider
$\phi_{j}(x_{j})_{0}=x_{j},\qquad\phi_{j}(x_{j})_{1}=1.$ (12)
Then every component of the global feature map is
$\phi(\vec{x})_{i_{1},\dots,i_{N}}=\prod_{j:\,i_{j}=0}x_{j},$ (13)
so the resulting monomials are square-free: each variable appears at most once. The induced kernel between two inputs is
$K(\vec{x},\vec{z})=\langle\phi(\vec{x}),\phi(\vec{z})\rangle=\prod_{j=1}^{N}(1+x_{j}z_{j}).$ (14)
Therefore the global feature map gives all possible square-free products between the components of the input vector, as a $2^{N}$-component tensor that expresses correlations in the vector. As we can see, in the standard way this would require an exponential amount of memory and time in the size of the input vector. Furthermore, if we want to apply a matrix to such a feature vector we would need a matrix $A$ of dimension $M\times 2^{N}$, which scales prohibitively.
However, by taking advantage of the properties of tensor networks we can use an MPO representation for such a matrix, of the form depicted in Fig. 4, provided that the output dimension $M$ is either small enough to be treated as a single physical output index or factorized as $M=\prod_{k}p_{k}$, so that $A$ can be represented with local input dimension $2$ and local output dimensions $p_{k}$. The advantage is lost if the required MPO bond dimensions grow exponentially. Under these conditions the operation can be performed through a sequence of local contractions. If the MPO has local input dimension $2$, local output dimensions $p_{k}$, and bond dimensions $\chi_{k}$, contracting it with the product feature map produces a compressed output with cores of size roughly $\chi_{k-1}\times p_{k}\times\chi_{k}$, at a cost polynomial in the local dimensions and bond dimensions. If the final $M$-dimensional output vector is materialized explicitly, an additional cost at least linear in $M$ is unavoidable. Thus the advantage comes from keeping the intermediate and, when possible, the output representation compressed.
@@
With this methodology we can apply an operation in which we weight elements of type $x_{1}x_{2}x_{3}x_{4}x_{5}$, $x_{1}x_{2}x_{3}x_{4}$, $x_{1}x_{2}x_{3}$, $x_{1}x_{2}$ and $x_{1}$, efficiently and with a reduced number of elements.
@@ 3.4 Data compression and denoising
In the field of big data there is a need to work with large volumes of data, which can be very expensive or unfeasible for certain systems. The tensor compression property that we have seen in representations such as MPS allow us to compress [12] and handle [49] large amounts of data. To do this we must first tensorize the data and then compress it. The tensorization step is not neutral: different reshapes, quantized tensor-train encodings, or variable orderings can lead to very different ranks. Denoising and compression are effective when the signal has low effective tensor rank and the discarded components are dominated by noise; for dense global correlations or structured noise, the required ranks may remain high.
@@ 3.5 Combinatorial optimization
Solving combinatorial optimization problems typically requires runtimes that grow rapidly with problem size. Quantum and quantum-inspired methods can provide useful heuristic and structural approaches for certain instances, especially when the cost function or the constraints admit compact tensor-network representations. However, for general NP-hard problems no polynomial-time exact method is known, and tensor-network ranks may still grow exponentially in the worst case. There are different types of combinatorial optimization solvers.
The first one is the one developed in [67], running in a black box. This algorithm starts from a cost tensor, $T$, whose subscripts are the components of the solution vector. This tensor will have the cost of all possible states, although it is not necessary to know the tensor explicitly: it is enough to have a function that gives us the cost as a function of its indexes. The method builds a TT approximation by tensor cross interpolation guided by the max-volume principle, that is, by selecting submatrices with large absolute determinant, or suitable rectangular generalizations, in order to obtain stable interpolatory approximations. This should not be interpreted simply as selecting singular values with the largest product. From this representation we will obtain, with high probability, the lowest cost element.
@@
A recent direction is to build feasibility directly into the tensor network. Feasible tensor networks and constrained MPS encode constraints in the support or block structure of the state, so that samples or amplitudes correspond only to configurations satisfying the prescribed constraints [45, 40]. This can reduce the burden of penalty tuning and avoids wasting probability mass on invalid configurations. The limitation is that constructing such networks may itself be problem-dependent, and the resulting ranks can still grow quickly for dense or global constraints.
@@
The third possibility is the use of imaginary-time evolution on the quantum state in uniform superposition of all possible combinations, equivalent to the application of a Hadamard gate to each qubit of a quantum circuit. Let $\Omega$ be the set of feasible configurations, let $C(s)$ be the cost function, and let
$D_{\beta}=\exp(-\beta C),\qquad|\psi_{0}\rangle=\frac{1}{\sqrt{|\Omega|}}\sum_{s\in\Omega}|s\rangle.$ (15)
where $D_{\beta}$ is the diagonal operator defined by $D_{\beta}|s\rangle=\exp(-\beta C(s))|s\rangle$. Then the normalized state
$|\psi_{\beta}\rangle=\frac{D_{\beta}|\psi_{0}\rangle}{\|D_{\beta}|\psi_{0}\rangle\|}$ (16)
has measurement probabilities proportional to $\exp(-2\beta C(s))$, so low-cost combinations are favored.
Imaginary-time evolution is not a deterministic closed-system unitary evolution. On quantum hardware it therefore requires additional mechanisms such as normalization, measurements, ancillas, postselection, or variational approximations. In classical tensor-network simulations, the non-unitary operator can be applied directly, provided that the resulting tensor-network ranks remain controlled [25]. If $S_{\star}$ is the set of minimizers, $C_{\star}$ is the minimum cost, and $\Delta$ is the gap to the best non-optimal value, then
$\Pr(s\notin S_{\star})\leq\frac{|\Omega|-|S_{\star}|}{|S_{\star}|}\exp(-2\beta\Delta).$ (17)
@@
Indeed, the total unnormalized weight of non-optimal states is at most $(|\Omega|-|S_{\star}|)\exp(-2\beta(C_{\star}+\Delta))$, whereas the optimal states contribute at least $|S_{\star}|\exp(-2\beta C_{\star})$, and dividing both bounds gives the claim. If we needed to take into account all possible combinations explicitly, we would need an unbearable amount of memory, resulting in a brute-force search. However, by combining the ability to compress information and simulate quantum states with the ability to apply non-unitary tensor-network operations, we can create such a state for various problems [3, 5, 4]. This can be done in an exact or approximate way, by recompressions, depending on how much memory we want to devote to it.
@@
The main limitation is that for NP-hard problems we may still need an exponential amount of memory and time to solve them exactly, because the tensor-network ranks can grow exponentially in the worst case. By making intelligent adjustments in the encoding and expression of the problem, the creation and contraction of the tensor network, and by applying various simplifications in the constraints, we can nevertheless obtain useful approximate or heuristic algorithms on structured instances.
@@
The fourth is the simulation of a physical system represented by a variational tensor network, with a cost function given as a Hamiltonian that we can write in MPS or MPO representation. In this case we will perform variational updates of the state tensor network, using techniques such as gradient descent, DMRG in the MPS setting, or related local sweep methods for other tensor-network geometries, in order to minimize the energy of the system and find the optimal combination. The effectiveness of these types of techniques depends very much on the problem to which they are applied and they can get stuck in local minima, especially for problems with constraints. In addition, there is the limitation that we must be able to express the cost function in a tensor network representation.
3.6 Software, contraction planning, and deployment
From an industrial perspective, the ansatz is only part of the engineering problem. Contraction-path optimization can be as important as the network architecture itself, because runtime and peak memory depend strongly on the chosen contraction tree, slicing strategy, and backend. Recent software ecosystems and standards, including ITensor/ITensors.jl workflows, quimb/cotengra-style contraction planning, TensorKrowch for machine-learning integration, GPU backends such as NVIDIA cuTensorNet, and community efforts toward standardized tensor-operation interfaces, make tensor-network prototyping and deployment considerably more practical than a few years ago [8, 43, 48]. For reproducible industrial evaluation, it is therefore important to report the tensorization, bond dimensions or truncation tolerances, contraction-path optimizer, hardware backend, numerical precision, peak memory, and wall-clock time, together with the accuracy metric relevant to the application.